OpenStudy (anonymous):

find the rule of : 5,15,45,......,3645

OpenStudy (anonymous):

I know that each number is multipled by 3, but i cannot figure out the rule.

OpenStudy (anonymous):

r=45/15=3 \[tn =ar ^{n-1}=5*3^{n-1},where r=3,a=5\]

OpenStudy (anonymous):

that's not it.. because if you put 100 in there it is only 300 and that is not correct..

OpenStudy (anonymous):

when n=100 \[tn=5*3^{100-1}=5*\left( 3 \right)^{99}=5*3*3*3........99\times ,which is \neq 300 \] but a very large no.

OpenStudy (anonymous):

\[3645=5*3^{n-1},3^{n-1}=\frac{ 3645 }{5 }=729=3^{6}\] n-1=6 n=6+1=7

OpenStudy (anonymous):

\[t3=5*3^{3-1}=5*9=45\] which is true.Because third term is 45.